# Photonic Doping of Epsilon-Near-Zero Bragg Microcavities

### v1

@@ 3.4 Magnetic Purcell Factor Enhancement in NC-Doped ENZ Cavity
The electric Purcell enhancement in NC-doped Bragg cavities can be investigated for different orders of NZI TM-polarized modes by exciting the structure with electric line currents placed at the electric hot spots of the modes. It should be noted that the use of electric or virtual magnetic line currents is purely a computational tool to excite the corresponding TM or TE modes; in practice, the TE- and TM-polarized modes can be excited by plane-wave illumination with the appropriate polarization. Figure 17 presents the corresponding calculations for the same cavity configuration as in Fig. 16, but with a reduced width of $W=12.5~\mu\mathrm{m}$ and with electric rather than magnetic line-current excitation. The resulting electric PF spectrum, shown in Fig. 17a, exhibits a series of sharp resonances corresponding to different orders of NZI TM-like modes. The field distributions for the lowest-order mode at $\lambda=1399.8~\mathrm{nm}$ are shown in Figs. 17b and 17c, illustrating the magnitudes of the electric field ($|E|$) and magnetic field ($|H|$), respectively. Both profiles demonstrate strong field localization within the NCs, with the inset enlargements highlighting the distinct electric hot spots characteristic of the resonant TM mode.
Figure 17: Electric Purcell enhancement in a Bragg cavity with $N=14$ cladding-layer pairs and width $W=12.5~\mu\mathrm{m}$, loaded with an array of 25 NCs of radius $R=150~\mathrm{nm}$ and period $p=500~\mathrm{nm}$. (a) Electric Purcell factor spectrum (b) Electric field distribution $|E|$ and (c) magnetic field distribution $|H|$ for the lowest-order TM-like mode at $\lambda=1399.8~\mathrm{nm}$. Insets highlight the strong field localization within the NCs and the distinct electric hot spots responsible for the observed enhancement.
4. Concluding Remarks
We have demonstrated that photonic doping of ENZ Bragg cavities with arrays of dielectric NPs provides a versatile platform for engineering high-$Q$ NZI modes with strong electric and magnetic responses. By combining the dispersive properties of near-cutoff Bragg cavities with the resonant character of embedded NPs, hybrid Bragg–Mie modes emerge that can be tuned not only through the conventional size parameter, but also via structural and external degrees of freedom such as cavity width, cutoff wavelength, and angle of incidence. These modes support both electric and magnetic hot spots with unprecedented confinement, leading to PF enhancements of more than two orders of magnitude beyond those of isolated NPs. Importantly, the approach is not limited to circular geometries, as demonstrated for square cross-section NPs, and extends naturally to higher-order cavity modes supported by sidewall reflections. The resulting isolated, spectrally pure resonances with dipolar or multipolar character realize optical-frequency analogues of PMC and EMNZ behavior, enabled by the emergence of magnetic and electric hot spots. Our findings establish NC-doped ENZ Bragg cavities as a powerful paradigm for achieving ultra-narrow-bandwidth resonances with extreme field localization, offering new opportunities for low-threshold nonlinear optics, magnetic dipole spectroscopy, and quantum light–matter interactions.
Appendix A
We consider the Mie coefficient corresponding to TE modes of an infinite lossless cylinder:
$a_{n}=\frac{x_{i}J_{n}(x_{i})J_{n}^{\prime}(x_{e})-x_{e}J_{n}(x_{e})J_{n}^{\prime}(x_{i})}{x_{i}J_{n}(x_{i}){H_{n}^{(1)}}^{\prime}(x_{e})-x_{e}H_{n}^{(1)}(x_{e})J_{n}^{\prime}(x_{i})}$ (18)
The extinction coefficient is proportional to the real part of $a_{n}$, which can be written in terms of the real nominator (N) and complex denominator (D) of the coefficient as:
$\Re(a_{n})\approx\Re\left(\frac{1}{D}\right)\cdot N=\frac{\Re(D)\cdot N}{\Re(D)^{2}+\Im(D)^{2}}$ (19)
where $\Re(D)$ and $\Im(D)$ are the real and imaginary parts of $D$. We show that in the ENZ regime, the real part of the Mie coefficient $a_{n}$ (or $b_{n}$) behaves like a Lorentzian function normalized to unity at resonance:
$L(\omega)=\frac{\gamma^{2}}{\Delta\omega^{2}+\gamma^{2}}.$ (20)
In this representation, $\Im(D)$ vanishes at the resonance frequency and thus representing the role of the resonance detuning $\Delta\omega$, while the $\Re(D)$ approaches zero as the surrounding refractive index tends to zero, corresponding to the narrowing Lorentzian linewidth $\gamma$.
First, for $n=0$, using the small-argument approximations:
$H_{0}^{(1)}(x_{e})$ $\approx\left(1-\frac{x_{e}^{2}}{4}\right)+i\cdot\frac{2}{\pi}\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right),$ (21)
$H_{0}^{(1)\prime}(x_{e})$ $\approx-\frac{x_{e}}{2}+i\cdot\left(-\frac{2}{\pi x_{e}}\right)$ (22)
the real and imaginary components of the denominator are obtained as:
$\Re(D)$ $=-\frac{x_{i}x_{e}}{2}J_{0}(x_{i})-x_{e}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right),$ (23)
$\Im(D)$ $=-\frac{2x_{i}J_{0}(x_{i})}{\pi x_{e}}-\frac{2x_{e}J_{0}^{\prime}(x_{i})}{\pi}\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right).$ (24)
And using the small-argument approximations for the Bessel function:
$J_{0}(x_{e})$ $\approx 1-\frac{x_{e}^{2}}{4},$ (25)
$J_{0}^{\prime}(x_{e})$ $\approx-\frac{x_{e}}{2},$ (26)
the simplified expression for the numerator becomes:
$N\approx-\frac{1}{2}x_{i}x_{e}J_{0}(x_{i})-x_{e}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right).$ (27)
Therefore, when $x_{e}\to 0$, the resonance condition is obtained from Eq. (24) as $J_{0}(x_{i})=0$, and for the resonance width we have:
$\gamma_{0}\propto-x_{e}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right),$ (28)
which tends to zero as $x_{e}\to 0$.
For $n>0$, using the small-argument approximations:
$H_{n}^{(1)}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n}-\frac{1}{(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+2}+i\cdot\frac{(n-1)!}{\pi}\left(\frac{2}{x_{e}}\right)^{n}$ (29)
${H_{n}^{(1)}}^{\prime}(x_{e})\approx\frac{n}{2n!}\left(\frac{x_{e}}{2}\right)^{n-1}-\frac{(n+2)}{2(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+1}-i\cdot\frac{n(n-1)!}{2\pi}\left(\frac{2}{x_{e}}\right)^{n+1}$ (30)
the real and imaginary parts of the denominator are given by:
$\Re(D)=$ $\frac{1}{2n!}\bigg[nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n-1}$ (31)
$\quad+\frac{1}{2(n+1)!}\bigg[-(n+2)x_{i}J_{n}(x_{i})+x_{o}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1}$
$\Im(D)=\frac{(n-1)!}{2\pi}\bigg[nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{2}{x_{e}}\right)^{n+1}$ (32)
Now we consider the numerator of the Mie coefficient $b_{n}$
$N(x_{e})=x_{i}J_{n}(x_{i})J_{n}^{\prime}(x_{e})-x_{e}J_{n}(x_{e})J_{n}^{\prime}(x_{i}),$
and expand the Bessel functions up to two terms:
$J_{n}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n}-\frac{1}{(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+2}$ (33)
$J_{n}^{\prime}(x_{e})\approx\frac{1}{2(n-1)!}\left(\frac{x_{e}}{2}\right)^{n-1}-\frac{1}{2(n+1)!}\left(\frac{x_{e}}{2}\right)^{n+1}$ (34)
to get:
$N$ $\approx\bigg[nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n-1}$ (35)
$-\frac{1}{2(n+1)!}\bigg[x_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1}.$
We see from Eq. 32 that the resonance condition is met by:
$nx_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})=0,$ (36)
which for $x_{e}\to 0$, reduces to:
$J_{n}(x_{i})=0.$ (37)
And for the resonance width we have:
$\gamma_{n}\propto\frac{-1}{2(n+1)!}\bigg[(n+2)x_{i}J_{n}(x_{i})-x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1},$ (38)
or:
$\gamma_{n}\propto\frac{1}{2(n+1)!}\bigg[x_{e}^{2}J_{n}^{\prime}(x_{i})\bigg]\left(\frac{x_{e}}{2}\right)^{n+1},$ (39)
which tends to zero as $x_{e}\to 0$.
In the case where the electric field is polarized parallel to the cylinder axis (TM polarization), the extinction efficiency is again proportional to the real part of the scattering coefficients $b_{n}$, expressed as:
Using the small-argument expansions for $n=0$:
$H_{0}^{(1)}(x_{e})$ $\approx\left(1-\frac{x_{e}^{2}}{4}\right)+i\cdot\frac{2}{\pi}\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right),$ (41)
$H_{0}^{(1)\prime}(x_{e})$ $\approx-\frac{x_{e}}{2}+i\cdot\left(-\frac{2}{\pi x_{e}}\right),$ (42)
$J_{0}(x_{e})$ $\approx 1-\frac{x_{e}^{2}}{4},$ (43)
$J_{0}^{\prime}(x_{e})$ $\approx-\frac{x_{e}}{2},$ (44)
the real and imaginary parts of the denominator become:
$\Re(D)$ $=-\frac{x_{e}^{2}}{2}J_{0}(x_{i})-x_{i}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right),$ (45)
$\Im(D)$ $=-\frac{2}{\pi}J_{0}(x_{i})-\frac{2x_{i}}{\pi}J_{0}^{\prime}(x_{i})\left(\ln\left(\frac{x_{e}}{2}\right)+\gamma\right)$ (46)
and the numerator becomes:
$N\approx-\frac{1}{2}x_{e}^{2}J_{0}(x_{i})-x_{i}J_{0}^{\prime}(x_{i})\left(1-\frac{x_{e}^{2}}{4}\right)$ (47)
From Eq. (45) and (46) we see that to remove the divergence of the denominator, we should have:
$J_{0}^{\prime}(x_{i})=0,$ (48)
and consequently, $\Re(D)$ vanishes when $x_{e}\to 0$. However, $\Im(D)$ remains non-vanishing in the small-size-parameter limit, preventing the emergence of a resonance condition. So, as it can also be inferred from Fig. 2b, the TM0 mode of the infinite cylinder does not support a true resonance.
For $n>0$ we use the small-argument approximations:
$H_{n}^{(1)}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n}+i\cdot\frac{(n-1)!}{\pi}\left(\frac{2}{x_{e}}\right)^{n}$ (49)
${H_{n}^{(1)}}^{\prime}(x_{e})\approx\frac{n}{2n!}\left(\frac{x_{e}}{2}\right)^{n-1}-i\cdot\frac{n(n-1)!}{2\pi}\left(\frac{2}{x_{e}}\right)^{n+1},$ (50)
to obtain the real and imaginary parts of the denominator as:
$\Re(D_{n})=\left(\frac{x_{e}}{2}\right)^{n}\cdot\frac{1}{n!}\left[\frac{n}{2}J_{n}(x_{i})-x_{i}J_{n}^{\prime}(x_{i})\right]+\left(\frac{x_{e}}{2}\right)^{n+2}\left[\frac{x_{i}J_{n}^{\prime}(x_{i})}{(n+1)!}-\frac{(n+2)}{2(n+1)!}J_{n}(x_{i})\right]$ (51)
$\Im(D_{n})=-\left(\frac{2}{x_{e}}\right)^{n}\cdot\frac{(n-1)!}{\pi}\left[\frac{n}{2}J_{n}(x_{i})+x_{i}J_{n}^{\prime}(x_{i})\right]$ (52)
Using only the leading-order terms of the Bessel functions:
$J_{n}(x_{e})\approx\frac{1}{n!}\left(\frac{x_{e}}{2}\right)^{n},\quad J_{n}^{\prime}(x_{e})\approx\frac{1}{2(n-1)!}\left(\frac{x_{e}}{2}\right)^{n-1}$ (53)
the numerator becomes:
$N\approx\left(\frac{x_{e}}{2}\right)^{n}\cdot\frac{1}{n!}\left[\frac{n}{2}J_{n}(x_{i})-x_{i}J_{n}^{\prime}(x_{i})\right]$ (54)
so, from Eqs. (32) and (33) we see that the resonance condition is satisfied when:
$\frac{n}{2}J_{n}(x_{i})+x_{i}J_{n}^{\prime}(x_{i})=0$ (55)
and for the resonance width we have:
$\gamma_{n}\propto\Re(D),$ (56)
indicating that the resonance width tends to zero as $x_{e}\to 0$.
Appendix B
Mie resonance of spherical NPs in ENZ media
To analyze the ENZ limit $x_{e}\to 0$, we use the small-argument expansions of the functions:
$\psi_{n}(x_{e})=\frac{x_{e}^{n+1}}{(2n+1)!!}+\mathcal{O}(x_{e}^{n+3}),\quad\psi_{n}^{\prime}(x_{e})=\frac{(n+1)x_{e}^{n}}{(2n+1)!!}+\mathcal{O}(x_{e}^{n+2})$ (57)
$\xi_{n}(x_{e})=\frac{x_{e}^{n+1}}{(2n+1)!!}+i\left(-\frac{(2n-1)!!}{x_{e}^{n}}\right)+\mathcal{O}(x_{e}^{n+3})+i\mathcal{O}(x_{e}^{2-n})$ (58)
$\xi_{n}^{\prime}(x_{e})=\frac{(n+1)x_{e}^{n}}{(2n+1)!!}+i\left(n\frac{(2n-1)!!}{x_{e}^{n+1}}\right)+\mathcal{O}(x_{e}^{n+2})+i\mathcal{O}(x_{e}^{1-n})$ (59)
and write the real part of $b_{n}$ in terms of its nominator ($N$) and denominator ($D$) as:
$\Re(a_{n})\approx\Re\left(\frac{1}{D}\right)\cdot N=\frac{\Re(D)\cdot N}{\Re(D)^{2}+\Im(D)^{2}}$ (60)
Then the numerator $N$, $\Re(D)$ and $\Im(D)$ are calculated as:
$N=\frac{x_{e}^{n+1}}{(2n+1)!!}\left((n+1)\psi_{n}(x_{i})-x_{i}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+3}),$ (61)
$\Re(D)=\frac{x_{e}^{n+1}}{(2n+1)!!}\left((n+1)\psi_{n}(x_{i})-x_{i}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+3}),$ (62)
$\Im(D)=\frac{(2n-1)!!}{x_{e}^{n}}\left(n\psi_{n}(x_{i})+x_{i}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{2-n}).$ (63)
From these results we can see that the dipolar or multipolar magnetic resonances occur when $\Im(D)=0$ or:
$n\psi_{n}(x_{i})+x_{i}\psi_{n}^{\prime}(x_{i})=0,$ (64)
which is consistent with the blue curves in Fig. 3b. For the resonance widths we get:
$\gamma_{n}\propto\frac{x_{e}^{n+1}}{(2n+1)!!}\left((n+1)\psi_{n}(x_{i})-x_{i}\psi_{n}^{\prime}(x_{i})\right)$ (65)
tending to zero as $x_{e}\to 0$.
Similarly, for the $a_{n}$ coefficients we obtain:
$N=\frac{x_{e}^{n}}{(2n+1)!!}\left((n+1)x_{i}\psi_{n}(x_{i})-x_{e}^{2}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+2}),$ (66)
$\Re(D)=\frac{x_{e}^{n}}{(2n+1)!!}\left((n+1)x_{i}\psi_{n}(x_{i})-x_{e}^{2}\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{n+2}),$ (67)
$\Im(D)=(2n-1)!!\left(\psi_{n}(x_{i})n\frac{x_{i}}{x_{e}^{n+1}}+\psi_{n}^{\prime}(x_{i})\right)+\mathcal{O}(x_{e}^{1-n}),$ (68)
indicating that in the ENZ regime ($x_{e}\to 0$), the resonance width, proportional to $\Re(D)$, approaches zero. Although the divergence of the first term in $\Im(D)$ suggests that resonances are suppressed in this regime, a more careful analysis which follows, reveals that for any arbitrarily small $x_{e}$, there always exist roots of $\Im(D)$ given by
$x_{i}=\frac{\alpha}{2}+\frac{1}{2n}\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}},$ (69)
located near the zeros $\alpha$ of the function $\psi_{n}(x)$, which is consistent with the red curves in Fig. 3b.
The roots of $\Im(D)$
Let $\alpha$ denote a simple zero of $\psi_{n}(x_{i})$, that is,
$\psi_{n}(\alpha)=0,\quad\text{and}\quad\psi_{n}^{\prime}(\alpha)\neq 0.$ (70)
Expanding $\psi_{n}(x_{i})$ near $\alpha$ gives
$\psi_{n}(x_{i})\approx\psi_{n}^{\prime}(\alpha)(x_{i}-\alpha),$ (71)
and also,
$\psi_{n}^{\prime}(x_{i})\approx\psi_{n}^{\prime}(\alpha).$ (72)
Substituting into the function
$f(x_{i})=\psi_{n}(x_{i})n\frac{x_{i}}{x_{e}^{n+1}}+\psi_{n}^{\prime}(x_{i}),$ (73)
we find
$f(x_{i})\approx\psi_{n}^{\prime}(\alpha)\left(n\frac{x_{i}(x_{i}-\alpha)}{x_{e}^{n+1}}+1\right).$ (74)
Setting $f(x_{i})=0$, we obtain
$nx_{i}(x_{i}-\alpha)+x_{e}^{n+1}=0,$ (75)
or equivalently,
$nx_{i}^{2}-n\alpha x_{i}+x_{e}^{n+1}=0.$ (76)
which has the solutions:
$x_{i}=\frac{\alpha}{2}\pm\frac{1}{2n}\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}}.$ (77)
As $x_{e}\to 0$, the square root behaves like
$\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}}\approx n\alpha,$ (78)
thus the ++ root approaches $x_{i}=\alpha$ while the - root approaches $x_{i}=0$. Since we seek a root near $\alpha$, we select the positive sign root:
$x_{i}=\frac{\alpha}{2}+\frac{1}{2n}\sqrt{n^{2}\alpha^{2}-4nx_{e}^{n+1}}.$} (79)
Appendix C:
.5 Electric and Magnetic Dipoles, Green Tensor, LDOS, and Purcell Factor
C.1. Conventions
We use time-harmonic fields with the phasor convention $e^{-i\omega t}$. Free-space constants are $\varepsilon_{0}$ and $\mu_{0}$, with $c=1/\sqrt{\varepsilon_{0}\mu_{0}}$ and $k_{0}=\omega/c$. Throughout, material permeabilities are taken as $\mu(\mathbf{r},\omega)=\mu_{0}$.
C.2. Electric dipole: field and Purcell factor
Applying the inhomogeneous vector Helmholtz equation
$\nabla\times\nabla\times\mathbf{E}(\mathbf{r})-k_{0}^{2}\,\varepsilon(\mathbf{r},\omega)\,\mathbf{E}(\mathbf{r})=i\omega\mu_{0}\,\mathbf{J}(\mathbf{r}),$ (80)
the electric dyadic Green tensor $\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)$ is defined as the solution of
$\nabla\times\nabla\times\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)-k_{0}^{2}\,\varepsilon(\mathbf{r},\omega)\,\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)=\mathbf{I}\,\delta(\mathbf{r}-\mathbf{r}^{\prime}),$ (81)
satisfying the outgoing-wave radiation condition. The total field produced by an arbitrary current distribution $\mathbf{J}(\mathbf{r}^{\prime})$ is obtained as
$\mathbf{E}(\mathbf{r})=i\omega\mu_{0}\int\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)\,\mathbf{J}(\mathbf{r}^{\prime})\,\mathrm{d}^{3}r^{\prime}.$ (82)
A point electric dipole $\mathbf{p}$ at $\mathbf{r}_{0}$ is represented by a polarization $\mathbf{P}(\mathbf{r})=\mathbf{p}\,\delta(\mathbf{r}-\mathbf{r}_{0})$, with bound current
$\mathbf{J}(\mathbf{r})=-\,i\omega\,\mathbf{P}(\mathbf{r})=-\,i\omega\,\mathbf{p}\,\delta(\mathbf{r}-\mathbf{r}_{0}).$ (83)
Substituting (83) into (82) gives
$\mathbf{E}(\mathbf{r})=\mu_{0}\omega^{2}\,\mathbf{G}(\mathbf{r},\mathbf{r}_{0};\omega)\,\mathbf{p},\qquad\mathbf{E}(\mathbf{r}_{0})=\mu_{0}\omega^{2}\,\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\,\mathbf{p}.$ (84)
We also write $\mathbf{p}=p_{0}\,\mathbf{u}_{p}$ with $|\mathbf{u}_{p}|=1$.
The normalized electric decay rate (Purcell factor) for a dipole oriented along $\mathbf{u}_{p}$ at $\mathbf{r}_{0}$ is
$\frac{\Gamma(\omega)}{\Gamma_{0}(\omega)}=\frac{6\pi}{k}\,\Im\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{p}\right\},\qquad k=\frac{\omega}{c}.$ (85)
From (84),
$\mathbf{u}_{p}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{p}=\frac{\mathbf{u}_{p}\!\cdot\!\mathbf{E}(\mathbf{r}_{0})}{\mu_{0}\omega^{2}\,p_{0}}.$ (86)
Choosing the (arbitrary) source phase $p_{0}=i|p_{0}|$ and using $\Im\{X/i\}=-\Re\{X\}$ for any complex scalar $X$, we obtain
$\Im\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{G}\!\cdot\!\mathbf{u}_{p}\right\}=-\,\frac{\Re\!\left\{\mathbf{u}_{p}\!\cdot\!\mathbf{E}(\mathbf{r}_{0})\right\}}{\mu_{0}\omega^{2}|p_{0}|}.$ (87)
Substituting (87) into (85) yields
C.3. Magnetic dipole: field and Purcell factor
We include a fictitious magnetic current density $\mathbf{M}$ to treat magnetic dipoles on the same footing as electric ones (phasor convention $e^{-i\omega t}$):
$\nabla\times\mathbf{E}$ $=i\omega\mu_{0}\,\mathbf{H}\;-\;\mathbf{M},$ (89)
$\nabla\times\mathbf{H}$ $=\mathbf{J}\;-\;i\omega\varepsilon_{0}\varepsilon(\mathbf{r},\omega)\,\mathbf{E}.$ (90)
Let $\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)$ be the electric dyadic Green tensor defined in (81). One may verify by direct substitution in (89)–(90) that the magnetic field generated by an arbitrary $\mathbf{M}$ is
$\mathbf{H}(\mathbf{r})=i\omega\,\varepsilon_{0}\varepsilon(\mathbf{r},\omega)\int\mathbf{G}(\mathbf{r},\mathbf{r}^{\prime};\omega)\,\mathbf{M}(\mathbf{r}^{\prime})\,\mathrm{d}^{3}r^{\prime}.$ (91)
A point magnetic dipole $\mathbf{m}=m_{0}\,\mathbf{u}_{m}$ with $|\mathbf{u}_{m}|=1$, located at $\mathbf{r}_{0}$ is represented by
$\mathbf{M}(\mathbf{r})=-\,i\omega\,\mathbf{m}\,\delta(\mathbf{r}-\mathbf{r}_{0}),$ (92)
so that, evaluating (91) at the source point and denoting the local (host) permittivity by $\varepsilon_{b}(\omega)=\varepsilon(\mathbf{r}_{0},\omega)$,
$\mathbf{H}(\mathbf{r})=\varepsilon_{0}\varepsilon_{b}\,\omega^{2}\,\mathbf{G}(\mathbf{r},\mathbf{r}_{0};\omega)\,\mathbf{m},\qquad\mathbf{H}(\mathbf{r}_{0})=\varepsilon_{0}\varepsilon_{b}\,\omega^{2}\,\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\,\mathbf{m}.$ (93)
Let $c_{b}=1/\sqrt{\varepsilon_{0}\varepsilon_{b}\mu_{0}}$ be the wave speed and $k_{b}=\omega/c_{b}$ the wavenumber in the host at $\mathbf{r}_{0}$. The magnetic Purcell factor (normalized decay rate) is
$\frac{\Gamma_{m}(\omega)}{\Gamma_{m,0}(\omega)}=\frac{6\pi}{k_{b}}\,\Im\!\left\{\mathbf{u}_{m}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{m}\right\}.$ (94)
Projecting (93) along $\mathbf{u}_{m}$ gives
$\mathbf{u}_{m}\!\cdot\!\mathbf{G}(\mathbf{r}_{0},\mathbf{r}_{0};\omega)\!\cdot\!\mathbf{u}_{m}=\frac{\mathbf{u}_{m}\!\cdot\!\mathbf{H}(\mathbf{r}_{0})}{\varepsilon_{0}\varepsilon_{b}\,\omega^{2}\,m_{0}}.$ (95)
Choosing the dipole phase $m_{0}=i|m_{0}|$ (so that $\Im\{X/i\}=-\Re\{X\}$) yields
[…]

### v2

@@ 3.4 Magnetic Purcell Factor Enhancement in NC-Doped ENZ Cavity
The electric Purcell enhancement in NC-doped Bragg cavities can be investigated for different orders of NZI TM-polarized modes by exciting the structure with electric line currents placed at the electric hot spots of the modes (see SI, Sec. D). It should be noted that the use of electric or virtual magnetic line currents is purely a computational tool to excite the corresponding TM or TE modes; in practice, the TE- and TM-polarized modes can be excited by plane-wave illumination with the appropriate polarization.
4 Conclusion
It is shown that photonic doping of ENZ Bragg cavities with periodic arrays of dielectric NCs provides a robust route to engineer high-$Q$ near-zero-index modes with well-controlled electric and magnetic character. On the analytical side, we demonstrated that embedding Mie resonators in an ENZ background generically suppresses radiative losses and narrows both electric and magnetic multipolar resonances for cylindrical and spherical inclusions. Numerically, we established that near-cutoff all-dielectric Bragg cavities doped with resonant silicon NC arrays support coupled Bragg–Mie modes with effective EMNZ response, quality factors approaching $10^{4}$, and magnetic-dipole Purcell enhancements exceeding $5\times 10^{3}$ for a $14\,\mu\mathrm{m}$-scale doped structure, far beyond what is achievable with isolated particles or undoped cavities.
The high-Q NZI modes can be selectively excited and tuned through structural and external parameters such as lattice period, cavity width, cutoff wavelength, and angle of incidence. This tunability enables access to spectrally isolated, predominantly single-multipole resonances and to controlled interference between electric- and magnetic-dipole-like channels. The high-$Q$ magnetic resonances of the NCs realize optical-frequency analogues of PMC and EMNZ behavior in a fully dielectric platform and provide large LDOS enhancements for both electric and magnetic emitters, with well-defined multipolar and polarization selectivity.
Beyond the specific configurations considered here, the concept of ENZ photonic doping is readily extendable to alternative dopant geometries, higher-order cavity modes, and three-dimensional architectures, as well as to active or nonlinear materials. We anticipate that NC-doped ENZ Bragg cavities will serve as a versatile building block for low-threshold and multipolar-selective lasing, magnetic-dipole and higher-order spectroscopy, enhanced magnetically driven nonlinear processes, and engineered quantum light–matter interactions.
Supporting Information
Analytical derivations of Mie resonances for cylindrical and spherical inclusions embedded in ENZ media, including the ENZ-limit behavior of TE and TM scattering coefficients and resonance linewidths. Formal definitions of electric and magnetic dipoles, dyadic Green tensors, local density of states, and Purcell factors are provided. Additional numerical results demonstrate electric Purcell enhancement in photonically-doped ENZ Bragg microcavities under TM-polarized excitation, together with corresponding near-field distributions. Supporting Information is available from the Wiley Online Library or from the author.
Acknowledgements
All authors acknowledge the support of the Flagship of Photonics Research and Innovation (PREIN) funded by the Research Council of Finland (grant no. 320165). JK also acknowledges the Magnus Ehrnrooth foundation for their PhD grant.